# smICA: an open source repository for mapping the concentration of fluorescenctly labeled molecules in living cells on the basis of confocal imaging combined with time-correlated single-photon counting

### v1

@@ Introduction
The fluorescently labeled mRNA (Cy5-m7GRNAegfp, stained with sulfo-Cy5 at the 5’ Cap) was synthesized and purified according to the previously published procedure12. The TRITC labeled dextran of molecular weights 4.4 kDa, 40 kDa and 155 kDa were bought from Sigma-Aldrich.
Fluorescence sorrelation spectroscopy - callibration, brighntess and cocnentration
@@
We apply the described methodology to the quantitative measurements of the concentration of the fluorescently labeled mRNA and its changes upon the time after injection. The cells were injected with CAP1-modified mRNA fluorescently labeled with the sulfo-Cy5 dye.12 The mRNA was encoding the green fluorescent protein. The initial concentration of injected mRNA was around 100 nM and varied from cell to cell. Concentration imaging was performed on three cells. Between the image acquisition, we performed FCS measurements to determine the concentration of mRNA molecules freely diffusing along the cytoplasm.
@@ Acknowledgments
TK: Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Visualization, Supervision, Writing - Original Draft, Writing - Review & Editing. GB: Investigation, Data Curation, Methodology, Validation, Writing - Original Draft, Writing - Review & Editing. JM: Investigation, Visualization, Data Curation, Writing - Review & Editing. KK: Investigation, Data Curation, Methodology, Writing - Original Draft, Writing - Review & Editing. MP: Writing - Review & Editing. AM: Resources. JJ: Resources, Funding acquisition. RH: Funding acquisition, Writing - Review & Editing.

### v2

# smICA: an open source repository for mapping the concentration of fluorescently labeled molecules in living cells on the basis of confocal imaging combined with time-correlated single-photon counting

@@ Introduction
The fluorescently labeled mRNA (Cy5-m7GRNAegfp, stained with sulfo-Cy5 at the 5’ Cap) was synthesized and purified according to the previously published procedure12. Briefly, mRNA was in vitro transcribed in the presence of azido-modified cap analog (N3-m7GpppG) and subsequently isolated with spin columns. After labeling with sulfo-Cy5-DBCO, the product was isolated with HPLC and characterized with gel electrophoresis. The TRITC labeled dextran of molecular weights 4.4 kDa, 40 kDa and 155 kDa were bought from Sigma-Aldrich.
Fluorescence correlation spectroscopy - callibration, brighntess and cocnentration
@@
We apply the described methodology to the quantitative measurements of the concentration of the fluorescently labeled mRNA and its changes upon the time after injection. The cells were injected with CAP-modified mRNA fluorescently labeled with the sulfo-Cy5 dye.12 The mRNA was encoding the green fluorescent protein. The initial concentration of injected mRNA was around 100 nM and varied from cell to cell. Concentration imaging was performed on three cells. Between the image acquisition, we performed FCS measurements to determine the concentration of mRNA molecules freely diffusing along the cytoplasm.
@@ Acknowledgments
TK: Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Visualization, Supervision, Writing - Original Draft, Writing - Review & Editing. GB: Investigation, Data Curation, Methodology, Validation, Writing - Original Draft, Writing - Review & Editing. JM: Investigation, Visualization, Data Curation, Writing - Review & Editing. KK: Investigation, Data Curation, Methodology, Software (tesitng), Writing - Original Draft, Writing - Review & Editing. MP: Methodology, Software (tesitng), Writing - Review & Editing. AM: Resources. JJ: Resources, Funding acquisition. RH: Funding acquisition, Writing - Review & Editing.